### v1

@@ Abstract
The origin of magnetic white dwarfs has been a long standing puzzle. Proposed origin mechanisms have included: fossil fields frozen in from the progenitor convective core; a dynamo in the progenitor envelope; crystallization dynamos in sufficiently cool white dwarfs; and accretion disk dynamos from white dwarf-white dwarf mergers or tidally shredded low mass stellar or planetary companions. Here we show how observational constraints on white dwarf magnetic field strengths, ages, and masses can be used to narrow down the viability of proposed origin mechanisms. Using the best available data, we find that the fossil field mechanism overpredicts the number of magnetic white dwarfs, which suggests, as supported by theoretical arguments, that the field is not actually frozen into the progenitor cores but diffuses before the white dwarfs forms. Crystallization dynamos, if operative, occur too late to explain the bulk of magnetic white dwarfs. And with progenitor envelope dynamos impeded by the theoretical challenge of depositing field from envelope to white dwarf core, the two disk dynamo mechanisms emerge as the field origin mechanisms most resilient to present constraints. The methods herein also reveal observational data gaps, and motivate future acquisition of more complete data.
@@
Magnetic fields provide an external signature of internal dynamics that helps to probe the evolution of stars and compact objects. For white dwarfs (WDs), approximately $400$ magnetic white dwarfs (MWDs) have been catalogued in the Montreal White Dwarf Database (MWDD), which contains about 30,000 objects [1] with magnetic fields ranging from $\sim 10^{3}$ to $\sim 10^{9}$ G.
Zeeman spectral line broadening and spectropolarimetry are the primary methods for measuring these WD magnetic fields [2, 3, 4]. The former provides a measure of the total magnetic field magnitude averaged over the surface of the WD $\langle|B|\rangle$ while spectropolarimetry provides information about the average value of the signed line of sight field over the WD surface $\langle B_{z}\rangle$. Unless both techniques are available, which is more likely for cooler WDs with fields in the range $1\lesssim\langle|B|\rangle\lesssim 100$MG [5], the values for reported magnetic fields are lower limits on the total field. If only the line of sight field is available, and the topology is assumed to be a dipole, then the total dipole field can be estimated, but this is still a lower limit on the total field.
@@
Complementarily, WDs undergo crystallization as they reach sufficiently low temperatures [9, 10]. As the WD solidifies into a crystalline lattice, heavier elements like ${}^{12}C$ and ${}^{16}O$ move towards the core, and lighter elements rise towards the outer layers, initiating a composition gradient [11]. The gravitational potential energy released is converted to the kinetic energy of buoyancy, that drive convection. This, along with WD rotation, may sustain a magnetic dynamo to amplify the magnetic field [12, 13]. Similar core crystallization has long been been thought to provide an important source of free energy for outer core dynamos in planetary bodies such as Earth [14]. But we must assess how many MWDs are old enough to be consistent with the timing of crystallization onset to be explained this way.
@@ Results
We next address the second condition for a viable fossil field, namely that the WD must form before the magnetic field decays. Large scale convective or turbulent dynamos require not only macroscopic transport but also a sufficiently high magnetic Reynolds number for the large-scale field stretching to overcome microphysical dissipation. Additionally, the turbulence must maintain a field growth transport term, such as kinetic helicity, that can overcome turbulent diffusion [e.g. 26]. But as convection wanes, the system will likely transition from a turbulent, dynamo-active state which is the most demanding, to turbulent but dynamo-inactive state, and eventually to laminar flow. Even a short time in the intermediate state of turbulence without a dynamo can be important because it means turbulent diffusion without compensating growth. We must compare therefore compare the turbulent diffusion time scale to the WD formation time scale. If the former is much shorter than the latter, then survival of the fossil field is highly questionable, as the field can decay before the WD fully forms.
@@
If residing below the line in Figure 1(b) were the only condition for a WD to retain a fossil field, that would vastly over-predict the number of magnetized WDs found in surveys. But the fact that the magnetic diffusion timescale is much shorter than the WD formation time scale plausibly explains why the fossil field mechanism is so inefficient, and likely not the dominant mechanism for WD magnetic field generation.
@@
Figure 3: Cooling curves for 3 distinct WD masses, with the onset of crystallization marked. a, WDs with field strengths below 50 MG. b, WDs with field strengths between 50–300 MG and 300–900 MG, showing their positions relative to the onset of crystallization.
@@
Overall, the data and plots reveal a notable increase in the fraction of magnetized WDs within the temperature range 5000-7500K, suggesting a potential link to the crystallization dynamo. This is further evidenced by the cooling curves in Figure 3, which show where the high field MWDs appear relative to the crystallization onset for different mass ranges. So despite the limited availability of data, as evidenced by the deficits indicated in Figure 2, crystallization cannot therefore be ruled out as an influence on magnetic field generation. However, given the presence of numerous magnetized WDs before the onset of crystallization it cannot be primary mechanism responsible for the magnetic fields of most MWDs.
@@
We assume a standard accretion disk model [43, 44] for which the mean fields and flows are axisymmetric, with turbulent transport supplied by the MRI [e.g. 45]. The Alfvén speed at each radius, $R$ for the total field in an MRI-saturated disk can be estimated from $\alpha_{ss}\beta\sim 0.1$ [46], where $\beta$ is the ratio of thermal to magnetic pressure, and $\alpha_{ss}$ is the Shakura-Sunyaev dimensionless parameter quantifying the turbulent stress that transports angular momentum. Typically, $0.0001\leq\alpha_{\text{ss}}\leq 0.1$ as measured in simulations or inferred from observations. Since the magnitude of the mean field dynamo-generated mean magnetic field in the disk is less than the total field magnitude, we have a relationship $\alpha_{ss}{\beta({\overline{v}}_{A}})\geq 0.1$ which implies that the squared Alfvén speed for the mean field is approximately ${\overline{v}_{A}}^{2}\leq 20\alpha_{ss}c_{\text{s}}^{2}$ [47], where the overline denotes quantities associated with the mean magnetic field. Here, $c_{\text{s}}$ is the sound speed which, from hydrostatic equilibrium, can also expressed as the product $c_{\text{s}}\simeq H\Omega_{\text{k}}$ where $H$ is the disk scale height, $\Omega_{\text{k}}$ is the Keplerian orbital frequency. From local mass conservation, the mass density of the in terms of the local accretion rate $\dot{M}$ is
@@
Figure 4: Comparison of Observed and Theoretical White Dwarf Magnetic Fields. The plot shows observed magnetic field strengths (y-axis) versus theoretical limits (x-axis) for magnetic white dwarfs. Dots and stars represent MWDs with different field strength ranges, with theoretical limits indicated in the legend. Solid and dashed lines denote the theoretical upper limits of poloidal and total magnetic field strength respectively, as derived in the text. The blue dashed line, overlaps exactly with the red solid line.
@@
As emphasized earlier, the observational methods often provide a lower limit to the total magnetic field strength estimated theoretically. In cases where the toroidal magnetic field is undetectable, the reported observed field strengths may more closely reflect the net signed poloidal field. If figure Accretion Dynamos are the most plausible source of most MWDs we therefore compare the observationally reported MWD fields to the maximum total WD magnetic fields predicted by pressure equilibrium (dotted lines) as well as the maximum poloidal fields (solid lines) associated with this equilibrium. To establish the pressure-limited magnetic field value, we use equations (2), (7) and $V_{\text{k}}=\sqrt{\frac{GM}{R}}$ at $R\sim R_{\text{WD}}$. We set $\dot{M}$ onto the WD to be a multiple of the Eddington rate of $\dot{M}_{\text{ed}}=L_{\text{ed}}/V_{\text{k}}^{2}=10^{-5}M_{\odot}\text{yr}^{-1}$ for $R_{\text{WD}}=10^{9}$cm, and $\alpha_{\text{ss}}=0.1$; any smaller $\alpha$ would produce even higher lines making this choice a conservative one. The dependence on $1/\alpha_{ss}^{1/2}$ for the lines corresponding to the total field arise from equation (2), since we choose a fixed accretion rate. The poloidal field lines are independent of $\alpha_{ss}$ as the poloidal field is reduced from the total by $\alpha_{ss}^{1/2}$.
The blue lines on the plot correspond to $\dot{M}=\dot{M}_{\text{ed}}$ and the red lines correspond to $\dot{M}=10\dot{M}_{\text{ed}}$. The dashed lines correpsond to the total field limit and the solid lines to the poloidal field. The multicolored middle line results because the blue dashed line overlaps exactly with the red solid line. Observed MWD fields are shown as dots and stars: blue dots indicate WDs with fields below the blue solid line; red dots indicate WDs with fields above the blue solid line but below the red solid line; and red stars represent WDs that lie below only the red dashed line. This analysis reveals that the estimated WD magnetic fields, derived from pressure equilibrium in an accretion scenario, can in principle explain the observed WD field strengths using plausible accretion rates $\dot{M}\leq 10{\dot{M}}_{\text{ed}}$.
@@ Discussion
Our analysis constrains the viable proposed mechanisms for MWDs when examined against existing observational and theoretical constraints. Accretion-driven dynamo mechanisms from tidally shredded companions or WD mergers are the hardest to presently rule out as an explanation for the magnetic fields in most MWDs, as summarized in Figure Accretion Dynamos are the most plausible source of most MWDs. More detailed theoretical investigation of accretion based models is particularly warranted, both to better understand the field generation in the disk and advection of the field onto the WD. And, as more data become available, the position of new WDs on updated versions of Figure Accretion Dynamos are the most plausible source of most MWDs will help to further constrain the plausibility of these mechanisms. In contrast, the hypothesis that magnetic fields in WDs are flux-frozen relics from progenitor stars seems unlikely, given that the progenitor turbulent diffusion timescales are much shorter than the WD formation timescale.
@@ Acknowledgements
We thank I. Caiazzo, J. Fuller, J. Nordhaus, L. Chamandy, and S. Ginzburg for related discussions. We acknowledge partial support from grants NSF PHY-2020249, and the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. We acknowledge support by the U.S DOE NNSA under Awards DE-NA0003856 and DE-NA0003842, DE-NA0004147, DE-NA0004144, and Subcontracts 536203 and 630138 with LANL and B632670 with LLNL through the Flash Center for Computational Science. Support from the U.S. DOE ARPA-E under Award DE-AR0001272 and U.S. DOE Office of Science, Fusion Energy Sciences under Award DE-SC0021990 is also acknowledged.

### v2

@@ Abstract
The origin of magnetic white dwarfs (MWDs) has been a long-standing puzzle. Proposed origin mechanisms have included: fossil fields frozen in from the progenitor convective core; a dynamo in the progenitor envelope; crystallization dynamos in sufficiently cool white dwarfs; and accretion disk dynamos from white dwarf–white dwarf mergers or tidally shredded low-mass stellar or planetary companions. Here we show how observational constraints on white dwarf magnetic field strengths, ages, and masses can be used to narrow down the viability of proposed origin mechanisms. Using data from both the Montreal White Dwarf Database and a 20 pc volume-limited sample from Gaia DR2, we find that the fossil field mechanism overpredicts the number of magnetic white dwarfs, which suggests, as supported by theoretical arguments, that the field is not actually frozen into the progenitor cores but diffuses before the white dwarf forms. Crystallization dynamos, occur too late to explain the bulk of magnetic white dwarfs. And with the progenitor envelope dynamos impeded by the theoretical challenge of depositing a field from envelope to white dwarf core, the two disk dynamo mechanisms emerge as the field origin mechanisms most resilient to present constraints, with mergers best able to explain the high mass, strongly magnetized young MWDs. The methods herein also reveal observational data gaps, and motivate future acquisition of more complete data.
@@
Magnetic fields provide an external signature of internal dynamics that helps to probe the evolution of stars and compact objects. For white dwarfs (WDs), approximately $400$ magnetic white dwarfs (MWDs) have been catalogued in the Montreal White Dwarf Database (MWDD), which contains about 30,000 objects [1] with magnetic fields ranging from $\sim 10^{3}$ to $\sim 10^{9}$ G. Zeeman splitting of Balmer lines and spectropolarimetry are the primary methods for measuring these WD magnetic fields [2, 3, 4, 5], with most MWD having been detected by the former [2], which provides a measure of the total magnetic field magnitude averaged over the surface of the WD $\langle|B|\rangle$. Spectropolarimetry provides information about the average value of the signed line of sight field over the WD surface $\langle B_{z}\rangle$ which would be a lower limit on the total field. If only the line of sight field is available, and the topology is assumed to be a dipole, then the total dipole field can be estimated as a lower limit on the total field. Both techniques are more likely to be available for cooler WDs with fields in the range $1\lesssim\langle|B|\rangle\lesssim 100$MG [6].
@@
Complementarily, WDs undergo crystallization as they reach sufficiently low temperatures [10, 11]. As the WD solidifies into a crystalline lattice, heavier elements like ${}^{16}O$ sink toward the center, while lighter elements like ${}^{12}C$ rise towards the outer layers, initiating a composition gradient [12]. The gravitational potential energy released is converted to the kinetic energy of buoyancy, that drive convection. This, along with WD rotation, may sustain a magnetic dynamo to amplify the magnetic field [13, 14]. Similar core crystallization has long been been thought to provide an important source of free energy for outer core dynamos in planetary bodies such as Earth [15]. But we must assess how many MWDs are old enough to be consistent with the timing of crystallization onset to be explained this way.
@@ Results
We next address the second condition for a viable fossil field, namely that the WD must form before the magnetic field decays. Large scale convective or turbulent dynamos require not only macroscopic transport but also a sufficiently high magnetic Reynolds number for the large-scale field stretching to overcome microphysical dissipation [27]. Additionally, the turbulence may require a field growth transport term, such as kinetic helicity, to overcome turbulent diffusion [e.g. 28]. But as convection wanes, the system will likely transition from a turbulent, dynamo-active state which is the most demanding, to turbulent but dynamo-inactive state that would quench the magnetic field by turbulent diffusion before evolving to laminar flow.
Even a short time in the intermediate state of turbulence without a dynamo can be important because it means turbulent diffusion without compensating growth. We must therefore compare the turbulent diffusion time scale to the WD formation time scale. If the former is much shorter than the latter, then survival of the fossil field is highly questionable, as the field can decay before the WD fully forms.
@@
As shown by [33] the detailed evolution of the convective boundary is likely complex, with the core convection zone shrinking during the main sequence phase, and the outer convective envelope later penetrating inward toward the core. However, this does not affect our argument above, as sustaining a large-scale dynamo still requires more stringent conditions than simply maintaining convection. Moreover, fields anchored in regions that do not ultimately become part of the white dwarf interior are not relevant to its magnetism. Although asteroseismology- based detections[34] suggest strong internal magnetic fields in the cores of red giants, these observations are made while the convective envelope is still present. The magnetic fields are inferred from their effects on stellar oscillation modes [35] in shells that may not correspond to regions that ultimately form the white dwarf. Our focus is on whether such fields can survive after the envelope is lost.
In addition to magnetic fields generated by convective core dynamos during the main sequence, some upper main sequence Ap/Bp stars, for example, exhibit strong, highly structured fields [36] already present at the zero-age Main sequence (ZAMS). These are thought to be inherited from earlier pre-main sequence stages of star formation, which can persist throughout the main sequence due to the presence of stable, non-convective envelopes. However, where these fields are anchored matters for their subsequent influence on WD fields. If these fields are anchored in the envelope, then they would be disconnected from the convective cores that become the WD and would not influence the WD fields when the envelope is lost. And if they are anchored in a sufficiently convective core, the turbulent diffusion would again wash away their signature and the fields would be subjected to whatever the turbulence and dynamo generate or decay, consistent with our previous discussion.
If residing below the line in Figure 1(b) were the only condition for a WD to retain a fossil field generated in the core, that would vastly over-predict the number of magnetized WDs found in surveys. But the fact that the magnetic diffusion timescale is much shorter than the WD formation time scale plausibly explains why the fossil field mechanism is so inefficient, and likely not the dominant mechanism for WD magnetic field generation.
@@
Figure 3: Cooling curves for 3 distinct WD masses, with the onset of crystallization marked. a, WDs with field strengths below 50 MG. b, WDs with field strengths between 50–300 MG and 300–900 MG, showing their positions relative to the onset of crystallization in the MWDD sample.
Figure 4: Cooling curves for 4 distinct WD masses, with the onset of crystallization marked. WDs with field strengths below 50 MG and 50–300 MG, showing their positions relative to the onset of crystallization in the volume-limited sample.
@@
Overall, the data and plots reveal a notable increase in the fraction of magnetized WDs within the temperature range 5000-7500K, suggesting a potential link to the crystallization dynamo. This is further evidenced by the cooling curves in Figures 3 and 4, which show where the high field MWDs appear relative to the crystallization onset for different mass ranges. So despite the limited availability of data, as evidenced by the deficits indicated in Figure 2, crystallization cannot therefore be ruled out as an influence on magnetic field generation.
To address potential biases in the full MWDD sample [8], we have also repeated our analysis using the volume-limited sample of 20 pc from Gaia DR 2 [42] which is a subset of the MWDD. This subset provides a more statistically balanced representation of WDs across a range of ages. We compare the results from this volume-limited sample, shown in figure 4 with those from the full MWDD sample to ensure that any data limitations affecting our conclusions are properly accounted for. Indeed the full sample over-represents younger WDs with strong magnetic fields, as seen from the fact that the subset shows that a higher fraction of its strongly magnetized WDs lie near or beyond the expected age of crystallization onset. However, given the presence of any magnetized WDs from any sample before the onset of crystallization demonstrates that this cannot be the only mechanism responsible for the magnetic fields of all MWDs.
In addition to the timing constraints, there remains the challenge that theoretical models [38][43] and simulations [44] of the the crystallization dynamo predict magnetic field strengths that are too weak to account for most observed MWDs.
@@
However, considerations arise when evaluating whether this model can explain the majority of the MWD population, specifically, the event frequency and the observed mass and age-distributions. Although a tidally shredded planetary companion can generate strong magnetic fields through accretion, it adds little to the mass of the final WD. As a result, MWDs produced through these mergers would require the progenitor star to already have been massive. WD-WD mergers, on the other hand, would naturally produce both strong magnetic fields and higher mass remnants. These could explain young, high-field WDs if the merger occurs during or close to the WD formation. However, whether such mergers occur frequently enough remains an open area of research: population synthesis models [51] provide estimates, though the accurate rates depend on mass ratios, binary evolution and orbital separations. Additional channels like wind-accelerated binary mergers [52] may help explain higher rates.
Magnetic fields in WDs appear to emerge at different stages in their evolution depending on mass. Strong magnetic fields are seen almost immediately after the formation of high mass WDs whereas in normal mass WDs, magnetism shows only after 1-2 Gyr of cooling. [8] We examined the dataset and find evidence for this trend, though the correlation is weak. If this trend survives more data, it further suggests that WD-WD mergers are the primary mechanism for the early appearance of strong magnetic fields in high mass WDs, while crystallization dynamos activated later in the cooling process are more common in normal-mass WDs. Accretion-based mechanisms might be able to explain some late emergence of magnetic fields if the onset of dynamo activity were delayed, for instance late stage tidal disruption events from a slow companion inspiral. These factors suggest that multiple field generation pathways contribute to the observed population.
@@
We assume a standard accretion disk model [54, 55] for which the mean fields and flows are axisymmetric, with turbulent transport supplied by the MRI [e.g. 56]. The Alfvén speed at each radius, $R$ for the total field in an MRI-saturated disk can be estimated from $\alpha_{ss}\beta\sim 0.1$ [57], where $\beta$ is the ratio of thermal to magnetic pressure, and $\alpha_{ss}$ is the Shakura-Sunyaev dimensionless parameter quantifying the turbulent stress that transports angular momentum. Typically, $0.0001\leq\alpha_{\text{ss}}\leq 0.1$ as measured in simulations or inferred from observations. Since the magnitude of the mean field dynamo-generated mean magnetic field in the disk is less than the total field magnitude, we have a relationship $\alpha_{ss}{\beta({\overline{v}}_{A}})\geq 0.1$ which implies that the squared Alfvén speed for the mean field is approximately ${\overline{v}_{A}}^{2}\leq 20\alpha_{ss}c_{\text{s}}^{2}$ [58], where the overline denotes quantities associated with the mean magnetic field. Here, $c_{\text{s}}$ is the sound speed which, from hydrostatic equilibrium, can also expressed as the product $c_{\text{s}}\simeq H\Omega_{\text{k}}$ where $H$ is the disk scale height, $\Omega_{\text{k}}$ is the Keplerian orbital frequency. From local mass conservation, the mass density of the in terms of the local accretion rate $\dot{M}$ is
@@
Figure 5: Comparison of Observed and Theoretical White Dwarf Magnetic Fields. The plot shows observed magnetic field strengths (y-axis) versus theoretical limits (x-axis) for magnetic white dwarfs. Dots and stars represent MWDs with different field strength ranges, with theoretical limits indicated in the legend. Solid and dashed lines denote the theoretical upper limits of poloidal and total magnetic field strength respectively, as derived in the text.
Figure 6: Comparison of Observed and Theoretical Magnetic Fields for a Volume-Limited (20pc) sample The plot shows observed magnetic field strengths (y-axis) versus theoretical limits (x-axis) for magnetic white dwarfs. Dots and stars represent MWDs with different field strength ranges, with theoretical limits indicated in the legend. Solid and dashed lines denote the theoretical upper limits of poloidal and total magnetic field strength respectively, as derived in the text.
@@
As emphasized earlier, the observational methods often provide a lower limit to the total magnetic field strength estimated theoretically. In cases where the toroidal magnetic field is undetectable, the reported observed field strengths may more closely reflect the net signed poloidal field. In figure Accretion Dynamos are the most plausible source of most MWDs we therefore compare the observationally reported MWD fields to the maximum total WD magnetic fields predicted by pressure equilibrium (dotted lines) as well as the maximum poloidal fields (solid lines) associated with this equilibrium. To establish the pressure-limited magnetic field value, we use equations (2), (7) and $V_{\text{k}}=\sqrt{\frac{GM}{R}}$ at $R\sim R_{\text{WD}}$. We set $\dot{M}$ onto the WD to be a multiple of the Eddington rate of $\dot{M}_{\text{ed}}=L_{\text{ed}}/V_{\text{k}}^{2}=10^{-5}M_{\odot}\text{yr}^{-1}$ for $R_{\text{WD}}=10^{9}$cm, and $\alpha_{\text{ss}}=0.1$; any smaller $\alpha$ would produce even higher lines making this choice a conservative one. The dependence on $1/\alpha_{ss}^{1/2}$ for the lines corresponding to the total field arise from equation (2), since we choose a fixed accretion rate. The poloidal field lines are independent of $\alpha_{ss}$ as the poloidal field is reduced from the total by $\alpha_{ss}^{1/2}$.
The blue lines on the plot correspond to $\dot{M}=\dot{M}_{\text{ed}}$ and the red lines correspond to $\dot{M}=10\dot{M}_{\text{ed}}$. The dashed lines correspond to the total field limit and the solid lines to the poloidal field. Observed MWD fields are shown as dots and stars: blue dots indicate WDs with fields below the blue solid line; red dots indicate WDs with fields above the blue solid line but below the red solid line; and red stars represent WDs that lie below only the red dashed line. This analysis reveals that the estimated WD magnetic fields, derived from pressure equilibrium in an accretion scenario, can in principle explain the observed WD field strengths using plausible accretion rates $\dot{M}\leq 10{\dot{M}}_{\text{ed}}$. To reduce the observational bias toward younger WDs present in the full MWDD sample as discussed earlier, we repeated this analysis with a 20pc volume-limited sample from Gaia DR2. As shown in Figure Accretion Dynamos are the most plausible source of most MWDs, the observed magnetic field strengths of WDs in this more complete sub-sample lie almost entirely within the poloidal field estimates derived from accretion at the Eddington rate. This supports the conclusion that the accretion model can explain the observed field strengths.
@@ Discussion
Our analysis constrains the viable proposed mechanisms for MWDs when examined against existing observational and theoretical constraints. Accretion-driven dynamo mechanisms from tidally shredded companions or WD mergers are the hardest to presently rule out as an explanation for the magnetic fields in most MWDs, as summarized in Figure Accretion Dynamos are the most plausible source of most MWDs. However, because the full sample may be biased toward younger, brighter, and more strongly magnetized white dwarfs due to detection thresholds, we also examine a less-biased, volume-limited sample from Gaia and find that the same trends persist, as shown in Figure Accretion Dynamos are the most plausible source of most MWDs. More detailed theoretical investigation of accretion based models is particularly warranted, both to better understand the field generation in the disk and advection of the field onto the WD. And, as more data becomes available, the position of new WDs on updated versions of Figure Accretion Dynamos are the most plausible source of most MWDs will help to further constrain the plausibility of these mechanisms. In contrast, the hypothesis that magnetic fields in WDs are flux-frozen relics from progenitor stars seems unlikely, given that the progenitor turbulent diffusion timescales are much shorter than the WD formation timescale.
@@ Acknowledgements
We thank J. Landstreet for thoughtful detailed comments that helped us to improve the manuscript. We thank I. Caiazzo, J. Fuller, J. Nordhaus, L. Chamandy, and S. Ginzburg for related discussions. We acknowledge partial support from the National Science Foundation (NSF), award no. PHY-2020249, and the Aspen Center for Physics, which is supported by the NSF, award no. PHY-2210452. We acknowledge support by the U.S DOE NNSA under award no. DE-NA0004144 and subcontracts no. 630138 and C4574 with Los Alamos National Laboratory to the Flash Center for Computational Science. Support from the U.S. DOE Office of Science, Fusion Energy Sciences under award no DE-SC0021990 is also acknowledged.